# Contracting with a Mechanism Designer

### v1

@@ 1.1 Our Contributions and Techniques
We observe that the PoA is approximately equal to the square of the PoDM. As discussed above, given the established connection between our model and the single-item auction setting, the first-best benchmark in PoA and the second-best benchmark in PoDM can be interpreted as the expected welfare and expected revenue, respectively. Meanwhile, the principal’s optimal expected utility corresponds to virtual value pricing, i.e., the virtual value of the price times the purchase probability. Since the expected revenue equals the expected virtual welfare, PoDM effectively measures the ratio between welfare and revenue in a modified instance where the buyer’s value distribution is replaced by the virtual value distribution from the original instance. Notably, this transformation does not preserve key structural properties: for example, a regular (resp. MHR) distribution may induce a virtual value distribution that is not regular (resp. MHR). Nevertheless, our analysis suggests that certain properties (e.g., probability tail bounds) implied by regularity or MHR conditions still (approximately) hold under this transformation.
@@
Intermediary with anonymous pricing.  We also consider an extended scenario in which the mechanism space of the intermediary is restricted to anonymous pricing mechanisms instead of all mechanisms and the agents have identical outcome distributions and cost distributions.  [^5] A natural question is that is there any monotonicity in the principal’s utility regarding two intermediaries with an inclusion relationship between their mechanism spaces? We give a negative answer to this question in Appendix B.
@@
Our result shows that the principal’s problem of designing an optimal contract remains a virtual value pricing with a single virtual agent whose value is equal to the first-order statistic among that of the real agents (see Theorem 6.1). We then quantify the intermediary’s impact measured by PoDM and PoA as we did in the basic scenario. Interestingly, we find that from the worst-case analysis perspective, restricting the intermediary’s mechanism space to pricing mechanisms has nearly the same impact on the principal’s utility as allowing unrestricted mechanisms (see the last two columns of Table 1). See Theorems 6.2 and 6.3 for details.
@@
Table 2: Bounds of PoDM and PoA for the unknown market size setting. Here $\kappa$ is the ratio of the smallest to the largest value in the support of the contribution distribution, and $\ell$ and $s$ are the largest and smallest value of the range of the number of agents, respectively.
Contract design for unknown the market size.  Considering that in practice it is difficult for principals to know the exact number of agents on the platform, we extend our model to a scenario in which the principal only knows that market size lies between $s$ and $\ell$.
Our results show that when the market size is unknown, both ratios suffer a multiplicative loss compared to the bounds established for known market sizes. Under regular distributions, the loss is a factor of $\frac{\ell}{s}$, while under MHR distributions, it reduces to $\mathcal{O}\left(\frac{\log\ell}{\log s}\right)$ (see Table 2). Interestingly, the former bound is obtained by designing the contract as if the market size were the largest possible, i.e., $\ell$, whereas the latter is achieved by designing the contract as if the market size is the smallest possible, i.e., $s$. Moreover, we provide an example illustrating that if the principal has no knowledge of the market size—i.e., if the number of agents ranges from 1 to infinity—then, regardless of the contract design, the gap between her utility and these benchmarks cannot be upper bounded. See Theorems 7.1, 7.2 and 7.3 for details.
@@ 1.2 Related Work
Intermediation in online advertising.  In online advertising markets, the ad exchange functions as a intermediary, as it has exclusive access to advertisers (i.e., buyers) seeking to purchase impressions (i.e., items) from the publisher (i.e., seller), conceptually aligning with the role of the intermediary in our model. Feldman et al. (2010) analyzes equilibria in a setting where advertisers compete for a unique impression from a central seller through intermediaries, with the flexibility to select which intermediary to engage with. Stavrogiannis et al. (2013, 2014) focus on advertisers who are not liquidity-constrained and bid through intermediaries but restrict attention to variants of the Vickrey auction. Balseiro and Candogan (2017) examines the optimal contract a intermediary offers to advertisers when their budgets and targeting criteria are private. Balseiro et al. (2021) explores multistage intermediaries in display advertising under second-price rules and characterizes a subgame perfect equilibrium among intermediaries and the seller. Zeithammer and Choi (2024) studies multi-layer auction markets under first-price auction rules. As noted in Section 1.1, when agents have identical success probabilities and cost distributions, the optimal contract design problem simplifies to a virtual value pricing problem, closely related to the publisher’s revenue optimization in online advertising. Our results on PoDM and PoA are also relevant, as they essentially compare publisher revenues in markets with and without ad exchanges.
@@
Bayesian auction design.  To bound the first-best and price of double marginalization in our model, we express the principal’s optimal utility through virtual value pricing, linking it to Bayesian auction design. Bayesian auction design focuses on revenue-optimal and efficiency-maximizing mechanisms under Bayesian priors on bidders’ valuations. Myerson (1981) introduces virtual values and characterizes the revenue-optimal auction as a virtual welfare maximizer. Aggarwal et al. (2009) examines the trade-off between revenue and efficiency, showing that optimal mechanisms often introduce inefficiencies. Several works explore simple yet near-optimal mechanisms: Dhangwatnotai et al. (2010) shows how a single value sample can approximate optimal pricing, while Jin et al. (2019b) analyzes the effectiveness of uniform pricing. Auctions with MHR distributions have been widely studied: Giannakopoulos and Zhu (2018), Jin et al. (2019a), and Giannakopoulos et al. (2021b) evaluates how simple mechanisms perform relative to optimal auctions under MHR assumptions. Yan (2011) studies the impact of value correlation on auction revenue, while Feng and Jin (2024) investigates when simple mechanisms can approximate optimal auctions even in irregular settings.
@@
Double marginalization and efficiency loss.  Double marginalization arises when upstream and downstream firms with market power set markups independently, leading to inflated prices and reduced welfare. Spengler (1950) shows that vertical integration can eliminate this inefficiency by internalizing markups—a structure conceptually similar to our benchmarks without an intermediary. Rey and Tirole (1986) further demonstrates that vertical contracts, such as two-part tariffs, can replicate the integrated outcome without requiring ownership consolidation. Perakis and Roels (2007) introduces the price of anarchy to quantify efficiency loss in decentralized supply chains under price-only contracts, a setting analogous to our anonymous pricing restriction. While their metric focuses on total system profit, our measures center on the principal’s utility. Cachon and Lariviere (2005) shows that revenue-sharing contracts can coordinate decentralized supply chains and flexibly divide surplus, though practical concerns such as administrative burden and incentive distortion may limit their effectiveness. Bernstein and Federgruen (2005) analyzes decentralized supply chains with competing retailers under demand uncertainty and demonstrates that price-discount sharing and buy-back contracts can restore efficiency, yet neither study quantifies inefficiency through formal ratios. More recent work by Choné et al. (2024) investigates vertical relationships under asymmetric information and strategic foreclosure, focusing on welfare impacts without defining explicit efficiency metrics. Loertscher and Marx (2022) further cautions that eliminating double marginalization through vertical integration may distort bargaining and harm competition. In contrast to these studies, we define and bound two formal efficiency measures—the price of anarchy (PoA) and the price of double marginalization (PoDM)—to quantify the intermediary’s impact from the principal’s perspective in environments with informational and contractual frictions.
@@ 2.1 The Three-Party Model
We consider a model that involves three parties: the principal, the intermediary, and $n$ agents. The principal has a task that can be completed by any one of the $n$ agents. However, the principal cannot directly access the agents. Instead, she can only interact with an intermediary who has access to the agents. Thus, our model can be described as a two-period interaction. At time $T=1$, the principal delegates the task to the intermediary via a contract: a mapping from outcomes to transfer amounts with the promise that if a particular outcome occurs, the principal will transfer the corresponding amount to the intermediary. At time $T=2$, the intermediary recruits an agent via a mechanism consisting of an allocation rule that maps the agents’ reported costs to the selected agent and a payment rule that specifies how much agents should be paid. The selected agent (if any) then completes the task at some cost. Details about the contract and mechanism designed by the principal and intermediary are discussed below.
We assume that the agents’ completion of the task can result in $m$ possible outcomes, where each outcome $j\in[m]$ is associated with a value $\lambda_{j}\geq 0$ representing the reward received by the principal if outcome $j$ occurs. Each agent $i\in[n]$ has a private cost $c_{i}$ for completing the task, independently drawn from a publicly known distribution $G_{i}$. If selected, agent $i$ induces a probability distribution $\bm{\rho}_{i}=(\rho_{i,1},\rho_{i,2},\ldots,\rho_{i,m})\in[0,1]^{m}$ over the outcome space $[m]$, where the probability of outcome $j$ is $\rho_{i,j}$ and $\sum_{j\in[m]}\rho_{i,j}=1$. The expected reward profile of agent $i$ is given by $\bm{\gamma}_{i}\triangleq(\lambda_{1}\cdot\rho_{i,1},\lambda_{2}\cdot\rho_{i,2},\ldots,\lambda_{m}\cdot\rho_{i,m})$, and the total expected reward generated by agent $i$ is $r_{i}\triangleq\sum_{j\in[m]}\lambda_{j}\cdot\rho_{i,j}$. Both the outcome distribution $\bm{\rho}_{i}$ and the expected reward $r_{i}$ are assumed to be public information for all parties. [^6] All our results hold even when the reward profile $\bm{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{m})$ is unobservable to the intermediary. See Appendix A for more explanation.
[…]
@@
Mechanism design between the intermediary and agents.  After the principal issues a contract $\bm{t}\in\mathbb{R}_{+}^{m}$ to the intermediary in time $T=1$, the intermediary chooses a mechanism $\mathcal{M}_{\bm{t}}=(\bm{x}_{\bm{t}},\bm{p}_{\bm{t}})$ consisting of the allocation rule $\bm{x}_{\bm{t}}$ and the payment rule $\bm{p}_{\bm{t}}$ in time $T=2$. Given any reported cost profile $\bm{c}=(c_{1},c_{2},\ldots,c_{n})$, the allocation rule $\bm{x}_{\bm{t}}\colon\mathbb{R}_{\geq 0}^{n}\to[0,1]^{n}$ maps the cost profile to the probability of agents being assigned the task. The allocation rule $\bm{x}_{\bm{t}}$ satisfies the feasibility constraint, i.e., $\sum_{i\in[n]}x_{\bm{t},i}(\bm{c})\leq 1$ for every reported cost profile $\bm{c}$. The payment rule $\bm{p}_{\bm{t}}\colon\mathbb{R}_{\geq 0}^{n}\to\mathbb{R}^{n}$ maps the cost profile to the payments from intermediary to each agent, i.e., $\bm{p}_{\bm{t}}=(p_{\bm{t},1},p_{\bm{t},2},\ldots,p_{\bm{t},n})$, where $p_{\bm{t},i}$ is the payment function to agent $i$. We assume the mechanism is dominant strategy incentive compatible (DSIC), i.e., every agent maximizes her utility by reporting her cost truthfully regardless of other agents’ reports, [^8] Due to the standard revelation principal (Myerson, 1981), this is without loss of generality. and ex post individual rational (ex-post IR), i.e., every agent receives non-negative utilities in the equilibrium regardless of other agents’ costs.
We assume risk-neutrality and quasi-linear utilities for all three parties. The principal, the intermediary, and agents aim to maximize their own utilities. The expected utility of the principal is $\sum_{i\in[n]}\left\langle\bm{1}-\bm{t},\bm{\gamma}_{i}\right\rangle\cdot x_{\bm{t},i}(\bm{c})$. The expected utilities of the intermediary and agent $i$ are $\sum_{i\in[n]}\left(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle\cdot x_{\bm{t},i}(\bm{c})-p_{\bm{t},i}(\bm{c})\right)$ and $p_{\bm{t},i}(\bm{c})-c_{i}\cdot x_{\bm{t},i}(\bm{c})$. [^9] Our model assumes that there is no moral hazard among agents, i.e., agents have no hidden action. A graphical illustration of this model is provided in Figure 1.
@@
Our three-party model can be viewed as the classic contract design in the (two-party) principal-agent models. Specifically, the intermediary (in the three-party model) can be considered as an agent (in the classic two-party contract design model), where all possible mechanisms forms her hidden action space.
@@
For an agent $i\in[n]$ with an expected reward $r_{i}$ and cost $c_{i}$, we define the following terms:

### v2

@@ 1.1 Our Contributions and Techniques
We observe that the PoA is approximately the square of the PoDM. As discussed above, given the established connection between our model and the single-item auction setting, the first-best benchmark in PoA and the second-best benchmark in PoDM can be interpreted as the expected welfare and expected revenue, respectively. Meanwhile, the principal’s optimal expected utility corresponds to virtual value pricing, i.e., the product of the virtual value of the price and the purchase probability. Since the expected revenue equals the expected virtual welfare, PoDM effectively measures the ratio between welfare and revenue in a modified instance where the buyer’s value distribution is replaced by the virtual value distribution from the original instance. Notably, this transformation does not preserve key structural properties of the original distribution: for example, a regular (resp., MHR) distribution may induce a virtual value distribution that is not regular (resp., MHR). Nevertheless, our analysis suggests that certain properties (e.g., probability tail bounds) implied by regularity or MHR conditions still hold approximately under this transformation.
@@
Intermediary with anonymous pricing. We also consider an extended scenario in which the intermediary’s mechanism space is restricted to anonymous pricing mechanisms rather than the full set of mechanisms, and the agents have identical outcome distributions and cost distributions.  [^5] A natural question is whether there is any monotonicity in the principal’s utility regarding two intermediaries with an inclusion relationship between their mechanism spaces? We give a negative answer to this question in Appendix B.
@@
Our result shows that the principal’s problem of designing an optimal contract remains a virtual value pricing involving a single virtual agent, whose value equals the first-order statistic of the real agents’ values (see Theorem 5.1). We then quantify the intermediary’s impact measured by PoDM and PoA as we did in the basic scenario. Interestingly, we find that in the worst-case analysis, restricting the intermediary’s mechanism space to pricing mechanisms has nearly the same impact on the principal’s utility as allowing unrestricted mechanisms (see the last two columns of Table 1). Formal statements are provided in Theorems 5.2 and 5.3.
@@
Table 2: Bounds of PoDM and PoA for the unknown market size setting. Here $\kappa$ is the ratio of the largest to the smallest value in the support of the contribution distribution, and $\ell$ and $s$ are the largest and smallest values of the range of the number of agents, respectively.
Contract design for the unknown market size. Considering that in practice it is difficult for principals to know the exact number of agents on the platform, we extend our model to a scenario in which the principal only knows that the market size lies in the interval $s,\ell]$.
Our results show that when the market size is unknown, both ratios suffer a multiplicative loss compared to the bounds established for known market sizes. Under regular distributions, the loss is a factor of $\frac{\ell}{s}$, while under MHR distributions, it reduces to $\mathcal{O}\left(\frac{\log\ell}{\log s}\right)$ (see Table 2). The former bound arises from optimizing the contract for any market size in $s:\ell]$, while the latter is uniquely attained by optimizing for the minimal size ss—optimizing for any other size fails to provide this guarantee. Moreover, we provide an example illustrating that if the principal has no knowledge of the market size—i.e., if the number of agents ranges from 1 to infinity—then, regardless of the contract design, the gap between her utility and these benchmarks cannot be upper-bounded. Formal statements are provided in Theorems 6.1, 6.2 and 6.3.
@@ 1.2 Related Work
Intermediation in online advertising. In online advertising markets, the ad exchange functions as an intermediary, as it has exclusive access to advertisers (i.e., buyers) seeking to purchase impressions (i.e., items) from the publisher (i.e., seller). This role conceptually aligns with the intermediary in our model. Feldman et al. (2010) analyze equilibrium in a setting where advertisers compete for a unique impression from a central seller through intermediaries, with the flexibility to choose which intermediary to engage. Stavrogiannis et al. (2013, 2014) focus on advertisers who are not liquidity-constrained and bid through intermediaries but restrict attention to variants of the Vickrey auction. Balseiro and Candogan (2017) examine the optimal contract an intermediary offers to advertisers when their budgets and targeting criteria are private. Balseiro et al. (2021) explore multistage intermediaries in display advertising under second-price rules and characterizes a subgame perfect equilibrium among intermediaries and the seller. Zeithammer and Choi (2024) study multi-layer auction markets under first-price auction rules. As noted in Section 1.1, when agents have identical success probabilities and cost distributions, the optimal contract design problem simplifies to a virtual value pricing problem, closely related to the publisher’s revenue optimization in online advertising. Our results on PoDM and PoA are also relevant, as they essentially compare publisher revenues in markets with and without ad exchanges.
@@
Bayesian auction design. To bound the price of double marginalization (PoDM) and price of anarchy (PoA) in our model, we express the principal’s optimal utility through virtual value pricing, thereby connecting it to the literature on Bayesian auction design, which studies revenue-optimal and efficiency-maximizing mechanisms under Bayesian priors over bidders’ valuations. Myerson (1981) introduces virtual values and characterizes the revenue-optimal auction as a virtual welfare maximizer. Aggarwal et al. (2009) examine the trade-off between revenue and efficiency, showing that optimal mechanisms often introduce inefficiencies. Several works explore simple yet near-optimal mechanisms: Dhangwatnotai et al. (2015) show how a single value sample can approximate optimal pricing, while Jin et al. (2019b) analyze the effectiveness of uniform pricing. Auctions with MHR distributions have been widely studied: Giannakopoulos and Zhu (2018), Jin et al. (2019a), and Giannakopoulos et al. (2021) evaluate how simple mechanisms perform relative to optimal auctions under MHR assumptions. Yan (2011) studies the impact of value correlation on auction revenue, while Feng and Jin (2024) investigate when simple mechanisms can approximate optimal auctions even in irregular settings.
@@
Double marginalization and efficiency loss. Double marginalization arises when upstream and downstream firms with market power set markups independently, leading to inflated prices and reduced welfare. Spengler (1950) shows that vertical integration can eliminate this inefficiency by internalizing markups—a structure conceptually similar to our benchmarks without an intermediary. Rey and Tirole (1986) further demonstrate that vertical contracts, such as two-part tariffs, can replicate the integrated outcome without requiring ownership consolidation. Perakis and Roels (2007) introduce the price of anarchy to quantify efficiency loss in decentralized supply chains under price-only contracts, a setting analogous to our anonymous pricing restriction. While their metric focuses on total system profit, our measures center on the principal’s utility. Cachon and Lariviere (2005) show that revenue-sharing contracts can coordinate decentralized supply chains and flexibly divide surplus, though practical concerns such as administrative burden and incentive distortion may limit their effectiveness. Bernstein and Federgruen (2005) analyze decentralized supply chains with competing retailers under demand uncertainty and demonstrate that price-discount sharing and buy-back contracts can restore efficiency, yet neither study quantifies inefficiency through formal ratios. More recent work by Choné et al. (2024) investigates vertical relationships under asymmetric information and strategic foreclosure, focusing on welfare impacts without defining explicit efficiency metrics. Loertscher and Marx (2022) further cautions that eliminating double marginalization through vertical integration may distort bargaining and harm competition. In contrast to these studies, we define and bound two formal efficiency measures—the price of anarchy (PoA) and the price of double marginalization (PoDM)—to quantify the intermediary’s impact from the principal’s perspective in environments with informational and contractual frictions.
@@ 2.1 The Three-Party Model
We consider a model involving three parties: a principal, an intermediary, and $n$ agents. The principal has a task that can be completed by any one of the agents. However, the principal cannot directly access the agents. Instead, she can only interact with an intermediary who has access to the agents. The interaction unfolds over two periods. At time $T=1$, the principal delegates the task to the intermediary via a contract: a mapping from outcomes to transfer amounts with the promise that if a particular outcome occurs, the principal will transfer the corresponding amount to the intermediary. At time $T=2$, the intermediary recruits an agent via a mechanism consisting of an allocation rule that maps the agents’ reported costs to the selected agent and a payment rule that specifies how much agents should be paid. The selected agent (if any) then completes the task at some cost. Details about the contract and mechanism designed by the principal and intermediary are discussed below.
We assume that the agents’ completion of the task can result in $m$ possible outcomes, where each outcome $j\in[m]$ is associated with a value $\lambda_{j}\geq 0$ representing the reward received by the principal if outcome $j$ occurs. Each agent $i\in[n]$ has a private cost $c_{i}$ for completing the task, independently drawn from a publicly known distribution $G_{i}$. If selected, agent $i$ induces a probability distribution $\bm{\rho}_{i}=(\rho_{i,1},\rho_{i,2},\ldots,\rho_{i,m})\in[0,1]^{m}$ over the outcome space $[m]$, where the probability of outcome $j$ is $\rho_{i,j}$ and $\sum_{j\in[m]}\rho_{i,j}=1$. The expected reward profile generated by agent $i$ is defined as $\bm{\gamma}_{i}\triangleq(\lambda_{1}\cdot\rho_{i,1},\lambda_{2}\cdot\rho_{i,2},\ldots,\lambda_{m}\cdot\rho_{i,m})$, and the total expected reward generated by agent $i$ is $r_{i}\triangleq\sum_{j\in[m]}\lambda_{j}\cdot\rho_{i,j}$. Both the outcome distribution $\bm{\rho}_{i}$ and the expected reward $r_{i}$ are assumed to be public information for all parties. [^6] All our results hold even when the reward profile $\bm{\lambda}=(\lambda_{1},\lambda_{2},\ldots,\lambda_{m})$ is unobservable to the intermediary. See Appendix A for more explanation.
Contract design between the principal and intermediary. The principal cannot directly observe the actions of the intermediary, that is, the mechanism selected by the intermediary or the agent subsequently chosen to complete the task. Instead, the principal observes only the stochastic outcome of the task resulting from this mechanism. [^7] This setting corresponds to the notion of moral hazard (hidden actions) in the classic principal-agent model. Consequently, the principal designs a contract that specifies the payment to the intermediary based on the observed outcome of the task. In our model, a contract is represented by a vector $\bm{t}=(t_{1},t_{2},\ldots,t_{m})\in\mathbb{R}_{+}^{m}$, where $t_{j}$ ($j\in[m]$) denotes the fraction of the reward transferred from the principal to the intermediary if outcome $j$ is realized. By requiring that transfers are always non-negative, we enforce the standard limited liability property in the contract theory. Note that $t_{j}$ may exceed $1$. As a sanity check, if agent $i$ completes the task, the expected payment from the principal to the intermediary is $\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle$, and the principal’s expected utility is $\left\langle\bm{1}-\bm{t},\bm{\gamma}_{i}\right\rangle$, where $\bm{1}=(1,1,\ldots,1)$ denotes the $m$-dimensional all-ones vector.
@@
Mechanism design between the intermediary and agents. After the principal issues a contract $\bm{t}\in\mathbb{R}_{+}^{m}$ to the intermediary in time $T=1$, the intermediary chooses a mechanism $\mathcal{M}_{\bm{t}}=(\bm{x}_{\bm{t}},\bm{p}_{\bm{t}})$ consisting of the allocation rule $\bm{x}_{\bm{t}}$ and the payment rule $\bm{p}_{\bm{t}}$ in time $T=2$. Given any reported cost profile $\bm{c}=(c_{1},c_{2},\ldots,c_{n})$, the allocation rule $\bm{x}_{\bm{t}}\colon\mathbb{R}_{\geq 0}^{n}\to[0,1]^{n}$ maps the cost profile to the probability of agents being assigned the task. The allocation rule $\bm{x}_{\bm{t}}$ satisfies the feasibility constraint, i.e., $\sum_{i\in[n]}x_{\bm{t},i}(\bm{c})\leq 1$ for every reported cost profile $\bm{c}$. The payment rule $\bm{p}_{\bm{t}}\colon\mathbb{R}_{\geq 0}^{n}\to\mathbb{R}^{n}$ maps the cost profile to the payments from intermediary to each agent, i.e., $\bm{p}_{\bm{t}}=(p_{\bm{t},1},p_{\bm{t},2},\ldots,p_{\bm{t},n})$, where $p_{\bm{t},i}$ is the payment function to agent $i$. We assume the mechanism is dominant strategy incentive compatible (DSIC), i.e., every agent maximizes her utility by reporting her cost truthfully regardless of other agents’ reports, [^8] Due to the standard revelation principle (Myerson, 1981), this is without loss of generality. and ex post individual rational (ex-post IR), i.e., every agent receives non-negative utilities in the equilibrium regardless of other agents’ costs.
We assume that all parties are risk-neutral and have quasi-linear utilities. The principal, the intermediary, and agents aim to maximize their own utilities. The expected utility of the principal is $\sum_{i\in[n]}\left\langle\bm{1}-\bm{t},\bm{\gamma}_{i}\right\rangle\cdot x_{\bm{t},i}(\bm{c})$. The expected utilities of the intermediary and agent $i$ are $\sum_{i\in[n]}\left(\left\langle\bm{t},\bm{\gamma}_{i}\right\rangle\cdot x_{\bm{t},i}(\bm{c})-p_{\bm{t},i}(\bm{c})\right)$ and $p_{\bm{t},i}(\bm{c})-c_{i}\cdot x_{\bm{t},i}(\bm{c})$. [^9] Our model assumes that there is no moral hazard among agents, i.e., agents have no hidden action. A graphical illustration of this model is provided in Figure 1.
@@
Our three-party model can be viewed as the classic contract design in the (two-party) principal-agent models. Specifically, the intermediary (in the three-party model) can be considered as an agent (in the classic two-party contract design model), where all possible mechanisms form her hidden action space.
@@
For an agent $i\in[n]$ with [^10] For the sake of brevity, we will henceforth abuse notation and refer to an agent with expected reward rather than expected reward generated by an agent. an expected reward $r_{i}$ and cost $c_{i}$, we define the following terms: